### v1

# Statistical closures from the Martin, Siggia and Rose approach to turbulence

Abstract
The goal of this paper is to study the statistical closures suggested by the Martin-Siggia and Rose approach to statistical turbulence. We find that the formalism leads to a Bethe-Salpeter equation for the three point correlation of the velocity field. In the leading order approximation this equation becomes an explicit expression, which reproduces the closure found in W D McComb and S R Yoffe, A formal derivation of the local energy transfer (LET) theory of homogeneous turbulence, J. Phys. A: Math. Theor. 50, 375501 (2017). This suggests ways to improve upon this closure by resumming the perturbative expansion and/or applying renormalization group techniques.
I Introduction
The goal of this paper is to discuss the solution of the closure problem in the field theory approach to statistical turbulence [1, 2, 3, 4], more specifcically within the Martin - Siggia and Rose (MSR) framework [5, 6, 7, 8, 9, 10]. We shall derive a Bethe-Salpeter equation [11, 12, 13, 14] for the three point correlation of the turbulent velocity. In the simplest (one-loop) approximation, this equation reduces to an explicit formula, which reproduces the closure advanced by McComb and Yoffe in ref. [15].
Let us describe the closure problem in greater detail. Our starting point is the randomly driven Navier-Stokes Equation (NSE) [16, 17, 18, 19, 20]
where is the molecular viscosity. We regard the pressure as a Legendre multiplier enforcing the constraint
In what follows, we shall assume the velocity fields are divergenceless and we shall omit the pressure term. The , which also obey , are the random forces, which we assume to be Gaussian with zero mean.
The problem is to compute the two-point correlation function
The usual approach is simply to multiply eq. (1) by and take the expectation value of the whole expression [16]. Then the right hand side may be computed from Novikov’ s formula [21, 22] (see below eq. (39)), but in the left hand side there is a term involving the three point function
We realize we have not found a closed equation but just the first rung in a hierarchy, since the corresponding equation for the three point function will involve a fourth order correlation and so on. The closure problem consists in finding a way to express the three point function eq. (4) as a functional of the correlation eq. (3) , thus truncating the hierarchy. Our goal is to see what truncation emerges from the MSR approach, and to compare it with existing results in the literature.
A key point in the analysis of ref. [15] is to formulate the problem in such a way that the full correlation function eq. (3) appears as an unknown, rather than performing a perturbative expansion for it. To replicate this within the MSR approach we must go beyond the 1-particle irreducible (1PI) effective action (EA) (see [23, 24, 25, 26, 27]). In this work we shall restrict ourselves to the 2-particle irreducible (2PI) EA [23, 28, 29, 30, 31]; for a discussion of 2PI techniques in a MSR framework see [32, 33]. We leave the consideration of higher irreducible EA’s for future work [34, 35, 36, 37, 38, 39, 40].
We shall show that the 2PIEA leads to a Bethe-Salpeter equation for the three point correlation, which in the simplest approximation, namely one-loop order, reduces to the closure derived in ref [15]. We believe this result validates the functional MSR approach to hidrodynamic turbulence, and opens up interesting new perspectives to be further investigated in future work.
The paper is organized as follows. In the next section II we present the 2PI MSR effective action; for a full discussion see. Then in section III we use it to derive a Bethe-Salpeter equation for the three point correlations, and in section IV we show that to leading order the Bethe-Salpeter equation reduces to McComb and Yoffe’s closure from ref. We conclude with some brief final remarks in section V.
II The Martin-Siggia and Rose 2PI Effective Action
Let us return to the randomly driven NSE eq. (1) and write the noise self correlation as
where
and
We shall assume the NSE (1) are the continuum limit of a discrete time evolution
where , ; is the projector on divergenceless fields
We also assume that the NSE admit only one solution for given initial conditions and realization of the driving forces, which moreover are regarded as independent of each other. This requires breaking random Galilean invariance [27]. For example, we may assume that the total linear momentum of the fluid vanishes, which in turn requires the random sources to average to zero over large scales. We shall assume the initial conditions have been set far enough in the past that any memory of them has been erased, and that a statistically homogeneous and isotropic configuration has been achieved.
We introduce a generating functional for the velocity field
Here, is the noise probability density function
We assume that normalization constants are already included in the integration measure. is the unique solution to the NSE for the given random driving. We may use the identity (cfr. eq. (1))
Under the discretization rules above the determinant has value [41]. We exponentiate the delta function adding a divergenceless auxiliary field and integrate over the driving forces to get
To obtain a more symmetric form we also add a source for the auxiliary field
Observe that
which shows that the expectation value of any product of auxiliary fields vanishes on shell, namely when the external sources vanish.
We are interested in a statistically homogeneous and isotropic situation where the mean velocity and mean auxiliary field vanish and the relevant information is carried by the correlation functions and (since ). To obtain self-consistent equations for the correlations we further introduce non local sources directly coupled to the products of fields
where is the Martin-Siggia-Rose (MSR) action
where
The first derivatives of the generating functional define the mean fields and correlations
where a double bracket such as means an irreducible correlation
When it is clear that the mean fields vanish we shall drop the double brackets.
The 2-particle irreducible effective action (2PIEA) is the full Legendre transform
whereby the mean fields obey the equations of motion
The 2PIEA has the structure [23]
where
is the sum of all 2 particle irreducible (2PI) vacuum bubbles built from the cubic vertex extracted from and irreducible correlations in the internal legs. Note that is independent of the mean fields, because the interaction is cubic.
Observe that a 2PI graph must be necessarily connected and that the simplest vacuum bubble built from cubic vertices contains two loops. In general, consider a vacuum bubble with vertices, loops and , and internal lines, where counts lines containing an correlation. Then we have
whereby
This shows there are no 2PI vacuum bubbles with just one loop.
III Bethe-Salpeter equation for three-point correlations
Let us write again the equation of motion derived from the variation of the 2PIEA with respect to in the case where all non local sources are set to zero,
where for any vector field
The correlations themselves are given by
We may use equations (34) and (36) to find the correlations on shell, where the mean fields and also . However, it is best to derive equations where the three point correlations appear explicitly. To this end, we take derivatives of equations (34) and (36) with respect to the local sources and in turn, and then set these sources to zero (thereby turning off the mean fields too). Recall the derivative rules
From the derivative of eq. (34)with respect to we get
Observe that the left hand side is just the expectation value . This may be contrasted to Novikov’s formula [21, 22]
so
On the other hand, let be any time such that . Then the solution of the NSE 1 for is uniquely determined by the velocity field at and the sources for , which are independent of the sources at . Therefore, we may write
In the limit the second derivative may be computed from the continuum limit of eq. (8)
no longer a random variable, and so we find that
also.
The derivative of eq. (34)with respect to yields
We finally must face the closure problem, that is, to express the three point functions in equations (38) and (44) in terms of two point correlations. Let us introduce the matrix of field products
so we may write eq. (36) as
Now the derivative with respect to yields
which may be cast as a Bethe-Salpeter equation
Similarly, the derivative with respect to yields
IV Derivation of the McComb-Yoffe closure
In the simplest approximation where we neglect we get
and similarly
Let us work out in detail the three point function
We introduce the Fourier transforms
Then
where
It is easily seen that
where
and
Since is antisymmetric in , we may discard this term and write
which is the closure found in [15].
V Final Remarks
In this work we have discussed the closure problem in statistical hydrodynamic turbulence within the framework of the Martin-Siggia and Rose 2PI effective action. The formalism leads to a Bethe-Salpeter equation for the three point functions which in the leading approximation reduces to an explicit closure, reproducing the results of ref [15].
We believe this result validates the use of functional methods and the MSR effective action in hydrodynamic turbulence. Over and above this, it immediately suggests strategies to go beyond this leading order results, using tools such as partial resummations [42, 43, 44, 45] and the renormalization group applied to higher irreducible effective actions [46, 47, 48, 49, 50, 51], which are already well developed in the literature.
Acknowledgements.
I thank the anonymous referee to Manuscript FW10261 ([27]) who suggested this project. E. C. acknowledges financial support from Universidad de Buenos Aires through Grant No. UBACYT 20020170100129BA, CONICET Grant No. PIP2017/19:11220170100817CO and ANPCyT Grant No. PICT 2018: 03684.References
- [1] G. Eyink and U. Frisch, Robert H. Kraichnan, in P. Davidson et al. (editors) A Voyage through Turbulence (Cambridge U.P., Cambridge, 2011).
- [2] Wyld, H.W., Jr. Formulation of the Theory of Turbulence in an Incompressible Fluid. Ann. Phys. 1961, 14, 143–165.
- [3] L. L. Lee, A Formulation of the Theory of Isotropic Hydromagnetic Turbulence in an Incompressible Fluid, Ann. Phys. 32, 292 (1965).
- [4] V. I. Belinicher and V.S. L’vov, A scale-invariant theory of fully developed hydrodynamic turbulence, Zh. Eksp. Teor. fiz. 93, 533 (1987) (Engl. Trans. Sov. Phys. JETP 66, 303 (1987)).
- [5] P.C. Martin, E.D. Siggia and H.A. Rose, Statistical Dynamics of Classical Systems, Phys. Rev. A8 423 (1973).
- [6] R Phythian, The functional formalism of classical statistical dynamics, J. Phys. A: Math. Gen., Vol. 10, No. 5, 777 (1977).
- [7] G.L. Eyink, Turbulence Noise, J. Stat. Phys. 83, 955 (1996).
- [8] A. Kamenev, Field Theory of Non-Equilibrium Systems, Cambridge University Press, Cambridge, U.K. (2011).
- [9] J. Zanella and E. Calzetta, Renormalization group and nonequilibrium action in stochastic field theory, Phys. Rev. E 66 036134 (2002).
- [10] C. de Dominicis and I. Giardina, Random fields and spin glasses: a field theory approach (Cambridge University Press, Cambridge, England, 2006).
- [11] Bethe, H.A. and Salpeter, E. E., A Relativistic Equation for Bound-State Problems, Phys. Rev. 84, 1232 (1951).
- [12] C. Itzykson and J-B. Zuber, Quantum field Theory, McGraw-Hill, New York (1980).
- [13] L. Tsang and J. Kong, Scattering of electromagnetic waves: Advanced topics, John Wiley, New York (2001).
- [14] M. Franco and E. Calzetta, Wave propagation in non-Gaussian random media, J. Phys. A: Math. Theor. 48 (2015) 045206.
- [15] W D McComb and S R Yoffe, A formal derivation of the local energy transfer (LET) theory of homogeneous turbulence, J. Phys. A: Math. Theor. 50, 375501 (2017).
- [16] Th. de Karman and L. Howarth, On the statistical theory of isotropic turbulence, Proc. R. Soc. A 164, 192 (1938).
- [17] T. von Kármán, Progress in the statistical theory of turbulence. Proc. Natl. Acad. Sci. USA 34, 530 (1948).
- [18] S. Chandrasekhar, The theory of turbulence, editado por E. Spiegel, Springer (2011).
- [19] A. S. Monin and A. M. Yaglom, Statistical Fluid Mechanics, MIT Press, 1971.
- [20] U. Frisch, Turbulence, the Legacy of A. N. Kolmogorov (Cambridge University Press, Cambridge, England, 1995).
- [21] E. A. Novikov, Random force method in turbulence theory, J. Exptl. Theoret. Phys. (U.S.S.R.) 44, 2159-2168 (1963) (Eng. trans. Sov. Phys. JETP 17, 1449 (1965)).
- [22] E. A. Novikov, Functionals and the random-force method in turbulence theory, J. Exptl. Theoret. Phys. (U.S.S.R.) 47, 1919-1926 (1964) (Eng. trans. Sov. Phys. JETP 20, 1290 (1965)).
- [23] E. Calzetta and B-L. Hu, Nonequilibrium Quantum Field Theory (Cambridge University Press, Cambridge, England, 2008).
- [24] J. Rammer, Quantum field theory of nonequilibrium states (Cambridge University Press, Cambridge (England), 2007)
- [25] F. M. Haehl, R. Loganayagam and M. Rangamani, Effective action for relativistic hydrodynamics: fluctuations, dissipation, and entropy inflow, JHEP 10, 194 (2018).
- [26] N. Mirón Granese, A. Kandus and E. Calzetta, Field Theory Approaches to Relativistic Hydrodynamics. Entropy 24, 1790 (2022).
- [27] E. Calzetta, Energy spectrum of non-Newtonian turbulence, Phys. Rev. Fluids 10, 054607 (2025).
- [28] M. E. Carrington, B. A. Meggison, and D. Pickering, 2PI effective action at four loop order in theory, Phys. Rev. D 94, 025018 (2016)
- [29] M.E. Carrington, G. Kunstatter, H. Zaraket, 2PI effective action and gauge dependence identities, Eur. Phys. J. C 42, 253–259 (2005).
- [30] E. Mottola, Gauge invariance in 2PI effective actions, Strong and Electroweak Matter 2002, pp. 432-436 (2003)
- [31] U. Reinosa, Renormalization and gauge symmetry for 2PI effective actions, in Strong and Electroweak Matter 2004, 316, World Scientific (2005) arXiv:hep-ph/0411255v1
- [32] E. Calzetta, Kadanoff-Baym equations for near-Kolmogorov turbulence, arXiv:0908.4068.
- [33] Tim Bode, The two-particle irreducible effective action for classical stochastic processes, J. Phys. A: Math. Theor. 55 (2022) 265401.
- [34] E. Calzetta and B. L. Hu, Decoherence of correlation histories, in Directions in General Relativity, Vol II: Brill Festschrift, edited by B. L. Hu and T. A. Jacobson Cambridge University Press, Cambridge, England, 1993.
- [35] E. Calzetta and B. L. Hu, Correlations, Decoherence, Dissipation, and Noise in Quantum Field Theory, in Heat Kernel Techniques and Quantum Gravity, edited by S. A. Fulling, Texas A&M Press, College Station, TX, 1995, hep-th/9501040.
- [36] E. Calzetta and B. L. Hu, Stochastic dynamics of correlations in quantum field theory: From the Schwinger-Dyson to Boltzmann-Langevin equation, Phys. Rev. D, VOLUME 61, 025012 (1999).
- [37] E. Calzetta, Fourth-order full quantum correlations from a Langevin–Schwinger–Dyson equation, J. Phys. A: Math. Theor. 42 (2009) 265401.
- [38] Jürgen Berges, n-particle irreducible effective action techniques for gauge theories, Phys. Rev. D, VOLUME 70, 105010 (2004).
- [39] M. E. Carrington, WeiJie Fu, T. Fugleberg, D. Pickering, and I. Russell, Bethe-Salpeter equations from the 4PI effective action, Phys. Rev. D 88, 085024 (2013).
- [40] M.E. Carrington, Techniques for calculations with nPI effective actions, EPJ Web of Conferences ,04013 (2015).
- [41] J. Zinn-Justin, Quantum Field Theory and Critical Phenomena (Clarendon Press, Oxford, 1993).
- [42] Victor L’vov and Itamar Procaccia, Exact resummations in the theory of hydrodynamic turbulence. I. The ball of locality and normal scaling, Phys. Rev. E52, 3840 (1995).
- [43] Victor L’vov and Itamar Procaccia, Exact resummations in the theory of hydrodynamic turbulence. II. A ladder to anomalous scaling, Phys. Rev. E52, 3858 (1995).
- [44] Victor L’vov and Itamar Procaccia, Exact resummations in the theory of hydrodynamic turbulence. III. Scenarios for anomalous scaling and intermittency, Phys. Rev. E53, 3468 (1995).
- [45] Victor L’vov and Itamar Procaccia, Towards a nonperturbative theory of hydrodynamic turbulence: Fusion rules, exact bridge relations, and anomalous viscous scaling functions, Phys. Rev. E54, 6268 (1996).
- [46] M. E. Carrington, Wei-Jie Fu, D. Pickering, and J.W. Pulver, Renormalization group methods and the 2PI effective action, Phys. Rev. D 91, 025003 (2015).
- [47] M. E. Carrington, S. A. Friesen, B. A. Meggison, C. D. Phillips, D. Pickering, and K. Sohrabi, 2PI effective theory at next-to-leading order using the functional renormalization group, Phys. Rev. D 97, 036005 (2018).
- [48] Margaret E. Carrington and Christopher D. Phillips, Four Loop Scalar Theory Using the Functional Renormalization Group, Universe 2019, 5, 9.
- [49] Federico Lamagna and Esteban Calzetta, A functional renormalization method for wave propagation in random media, J. Phys. A: Math. Theor. 50 (2017) 315102.
- [50] Jean-Paul Blaizot, Jan M. Pawlowski, Urko Reinosa, Functional renormalization group and 2PI effective action formalism, Annals of Physics 431 (2021) 168549.
- [51] Peter Millington and Paul M Saffin, Vertex functions and their flow equations from the 2PI effective action, J. Phys. A: Math. Theor. 55 (2022) 435402.